### v2

# Minibatch Optimal Transport and Perplexity Bound Estimation in Discrete Flow Matching

Abstract
Outperforming autoregressive models on categorical data distributions, such as textual data, remains challenging for continuous diffusion and flow models. Discrete flow matching, a recent framework for modeling categorical data, has shown competitive performance with autoregressive models. Despite its similarities with continuous flow matching, the rectification strategy applied in the continuous version does not directly extend to the discrete one due to the inherent stochasticity of discrete paths. This limitation necessitates exploring alternative methods to minimize state transitions during generation. To address this, we propose a dynamic-optimal-transport-like minimization objective for discrete flows with convex interpolants and derive its equivalent Kantorovich formulation. The latter defines transport cost solely in terms of inter-state similarity and is optimized using a minibatch strategy. Another limitation we address in the discrete flow framework is model evaluation. Unlike continuous flows, wherein the instantaneous change of variables enables density estimation, discrete models lack a similar mechanism due to the inherent non-determinism and discontinuity of their paths. To alleviate this issue, we propose an upper bound on the perplexity of discrete flow models, enabling performance evaluation and comparison with other methods.
Machine Learning, ICML
1 Introduction
Modeling distributions of data is a central task in machine learning. In the context of data distributions with continuous support, recent advancements of continuous diffusion and flow models (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2020; Lipman et al., 2023) have demonstrated impressive capabilities in generating data samples and performing density estimation (Song et al., 2021a, b; Chen et al., 2018) . Rectified flows, in particular, have gained attention recently as they enable high quality data generation with very few integration steps. Unfortunately, the application of continuous diffusion and flow models to categorical data distributions remains limited, as they generally underperform compared to autoregressive models in this setting (Chen et al., 2023; Gulrajani & Hashimoto, 2024; Li et al., 2022; Dieleman et al., 2022; Strudel et al., 2022).
To address this, recent research has focused on the development of discrete diffusion (Austin et al., 2021; Campbell et al., 2022; Meng et al., 2022; Lou et al., 2024; Sahoo et al., 2024; Ou et al., 2024; Shi et al., 2024) and discrete flow models (Campbell et al., 2024; Gat et al., 2024), which intuitively should be more suitable to model categorical data. In addition, they enable infilling and could potentially speed up generation. In this work, we focus on the discrete flow framework, as this approach increases the design space of possible probability evolutions as compared to discrete diffusion, by enabling the usage of arbitrary couplings and inner dynamics. In addition, discrete and continuous flows share many similarities, facilitating the adoption of results from continuous flow matching to the discrete case. However, despite similarities between the two frameworks, some fundamental differences exist. Such differences are mostly the result of the nonexistence of a discrete flows formulation, in which sample paths are deterministic.
A major implication of non-deterministic sample paths is that we can no longer use the rectification strategy presented in Liu et al. (2022). As paths in the discrete flow framework are ‘sequences’ of states, in this paper, we explore other ways to minimize the number of jumps/changes between states, which can be seen as the discrete version of minimizing path length. More generally, using similarity metrics between states, we attempt to find ways of defining flows that minimize the number of jumps weighted by the dissimilarity between states. This leads to defining a minimization problem in the context of discrete flow matching that resembles the dynamic formulation of optimal transport. More importantly, we then derive its equivalent Kantorovich formulation, where the cost function depends explicitly only on the similarity metric between states, and where the transport cost can be optimized using the minibatch strategy of Tong et al. (2023); Fatras et al. (2021). This result can be seen as a categorical-distribution version of the Benamou-Brenier theorem when the conditional flows are formulated as convex interpolants, which themselves can be seen as the categorical equivalent of the shortest paths conditional flows in the continuous framework. Interestingly, when the similarity metric in the dynamic formulation is chosen to be the discrete metric, the cost function in the Kantorovich formulation turns out to be the Hamming distance. Furthermore, when the similarity metric is the $L_{2}$ norm, the cost function is also the $L_{2}$ norm, as in the continuous case.
This contribution allows the reduction of the number of transitions in the generative process, but it also enables one to make better use of the similarity between the source and data distribution. Therefore it is feasible to construct a chain of flows, where each flow treats the previous modeled distribution as the source distribution, and refines it in order to make it more similar to the data distribution.
An additional implication of stochastic crossing paths in discrete flows is that one cannot use an equivalent of the instantaneous change of variable formula (Chen et al., 2018) for estimating the probability of a point being sampled by the model. This means that other approaches should be devised for estimating the perplexity. Inspired by the bound provided in Lou et al. (2024), we derive an upper bound of the perplexity in the discrete flow matching framework. This bound enables us to evaluate the performance of our model and compare it to other methods.
In summary, the main contributions of this paper include:
  • Defining a dynamic-OT-like objective to minimize the expected number of jumps between dissimilar states during generation, and deriving the equivalent Kantorovich formulation for convex interpolant conditional flows.
  • Deriving a perplexity upper bound in the framework of discrete flow models that enables evaluation and comparison with other approaches.
  • Proposing to construct a chain of flows which could particularly benefit from minibatch OT training.
2 Preliminaries and Notation
2.1 Discrete-Time Markov Chains Over Finite-State Spaces
A discrete-time Markov Chain in a finite-state space is a stochastic process $X_{1},X_{2},\ldots,X_{T}$, where each state $X_{t}$ depends solely on the preceding state $X_{t-1}$. The states $X_{t}$ can take on any value from the set $\{1,2,\ldots,S\}$, where $S$ denotes the total number of possible states, and $T$ represents the number of time steps. The probability of being in state $x$ at time $t$ is
$p_{t}(X_{t}=x)=\sum_{y=1}^{S}p(X_{t}=x,X_{t-1}=y)=\sum_{y=1}^{S}p_{t|t-1}(X_{t}=x|X_{t-1}=y)p_{t-1}(X_{t-1}=y).$ (1)
If we place all such probabilities $p_{t}(X_{t}=x)$ in a vector $s_{t}$ of shape $S\times 1$, such that $s_{t}(x)=p_{t}(X_{t}=x)$, then from above we can deduce that
$s_{t}=Ps_{t-1},$ (2)
where $P(x,y)=p_{t|t-1}(X_{t}=x|X_{t-1}=y)$. Given an initial probability distribution $s_{0}$ over states, the equation above fully determines the evolution of the probability over states with respect to time.
2.2 Continuous-Time Markov Chains Over Finite-State Spaces (Discrete Diffusion)
It is possible to define a stochastic process with the Markov property in finite-state spaces, for $t\in[0,T]$, (Anderson, 2012). As previously, we can define a discrete-time process, on time points $\{0,\epsilon,...,T-\epsilon,T\}$, such that there is $\epsilon$ probability of activating the previous transition mechanism when progressing from time $t-\epsilon$ to $t$, otherwise we stay where we are with probability ($1-\epsilon$). Removing the random variables to simplify notation, we have
$p_{t}(x)=(1-\epsilon)p_{t-\epsilon}(x)+\epsilon\sum_{y=1}^{S}p_{t|t-\epsilon}(x|y)p_{t-\epsilon}(y).$ (3)
We notice that when $\epsilon=1$ the equation above coincides with Equation (1), and in addition as before we can write Equation (3) in matrix form
$s_{t}=(1-\epsilon)s_{t-\epsilon}+\epsilon Ps_{t-\epsilon}=\left(I+\epsilon(P-I)\right)s_{t-\epsilon}=\left(I+\epsilon Q\right)s_{t-\epsilon}\text{ , where }Q=P-I.$ (4)
From Equation (4), we see that $\frac{s_{t}-s_{t-\epsilon}}{\epsilon}=Qs_{t-\epsilon}$, which when taking the limit $\epsilon\rightarrow 0$ becomes $\frac{ds_{t}}{dt}=Qs_{t}$. Given an initial probability distribution $s_{0}$ over states, the equation above fully determines the evolution (flow) of the probability $p_{t}$ over states with respect to time. Indeed, the distribution over states at time $t$ is $s_{t}=e^{tQ}s_{0}$. This formulation can be generalized, such that $Q$ is allowed to evolve with time,
$\frac{ds_{t}}{dt}=Q_{t}s_{t}.$ (5)
For the choice $Q_{t}=\sigma^{{}^{\prime}}(t)Q$, where $\sigma$ is monotonically increasing, $\sigma(0)=0$ and $\lim_{t\rightarrow 1}\sigma(t)=T$, we have $s_{t}=e^{\sigma(t)Q}s_{0}$. Matrices $Q$ must satisfy the properties of transition-rate matrices (Suhov & Kelbert, 2008), that is, they have non-negative non-diagonal entries, and the elements in each column add to 00. The choice for $Q$ is made such that: a) the matrix exponential $e^{\sigma(t)Q}$ is easy to calculate and $\boldsymbol{z}_{1}$ is an easy reference distribution to sample from (Austin et al., 2021; Campbell et al., 2022). Unfortunately, these conditions greatly restrict the design space in this framework.
2.3 Discrete Flow Matching
In order to expand the design space of discrete diffusion models, Campbell et al. (2024); Gat et al. (2024) introduce discrete flow matching. In this paper we will focus on the approach and use the notation presented in Gat et al. (2024). In discrete sequence modeling, a sequence $x$ is denoted as an array of $L$ elements $(x^{1},x^{2},\dots,x^{L})$. Each position $i$, within this sequence is filled with an element $x^{i}$ from a vocabulary of size $V$. Consequently, the entire set of possible sequences is given by $\mathcal{D}=[V]^{L}$, where $[V]=\{1,\dots,V\}$.
We denote with $p^{i}(x^{i})$ the marginal of $p$ at position $i$, i.e., $p(x^{i})=\sum_{x^{-i}}p(x)$, where $x^{-i}=(x^{1}\dots,x^{i-1},x^{i+1},\dots x^{L})$. The following delta function notation will be particularly useful,
$\delta_{y}(x)=\prod_{i=1}^{N}\delta_{y^{i}}(x^{i}),\text{ where }\delta_{y^{i}}(x^{i})=\begin{cases}1&\text{if }x^{i}=y^{i}\\ 0&\text{if }x^{i}\neq y^{i}\end{cases}.$ (6)
2.3.1 Probability Flows and Velocities
In discrete flow matching (Gat et al., 2024), the goal is to acquire a flow $p_{t}:[0,1]\times[V]^{L}\rightarrow[0,1]$ constrained by $\sum_{z\in[V]^{L}}p_{t}(z)=1$ that transforms source (reference) distributions $X_{0}\sim p$ to target (data) distributions $X_{1}\sim q$. The flow is completely defined by the choice of a probability velocity $u_{t}:[0,1]\times[V]^{L}\rightarrow\mathbb{R}^{L\times V}$, such that $u_{t}(z)=(u^{1}_{t}(z),\dots,u^{i}_{t}(z),\dots,u^{L}_{t}(z))$ and $u^{i}_{t}:[0,1]\times[V]^{L}\rightarrow\mathbb{R}^{V}$, where $u^{i}_{t}(z)[x^{i}\neq z^{i}]\geq 0$ and $\sum_{x^{i}\in[V]}u^{i}_{t}(z)[x^{i}]=0$, for each $i$. The update rule of the probability over states when going from time $t$ to $t+\epsilon$ is defined independently for each position in the sequence as follows $p^{i}_{t+\epsilon|t}(x^{i}|x_{t})=\delta_{x^{i}_{t}}(x^{i})+u^{i}_{t}(x^{i},x_{t})$, where we used $u^{i}_{t}(x^{i},z):=u^{i}_{t}(z)[x^{i}]$. Therefore, we can see that as in the framework of Markov chains, the probability over the states in the next step depends solely on the current state, and that $u_{t}$ plays a similar role to a transition-rate matrix $\bar{Q}_{t}$, completely determining the flow. As such, if we learn to approximate the probability velocity $u_{t}(z)$ using a neural network $u_{t}(z;\theta):[0,1]\times[V]^{L}\rightarrow\mathbb{R}^{L\times V}$, we can sample from $p$ and generate data from $q$, using the previous update rule. Before modeling the probability velocity $u_{t}(z)$ however, one must first design an appropriate flow $p_{t}(z)$ that has a suitable, practically learnable corresponding $u_{t}(z)$.
2.3.2 Conditional Probability Flows
Since at time $t=0$ and $t=1$ we must have $p_{0}=p$ and $p_{1}=q$ respectively, we are already restricted regarding the endpoints of the flow. A trivial way to satisfy such constraints is to define
$p_{t}(x)=\sum_{x_{0},x_{1}\in\mathcal{D}}p_{t}(x|x_{0},x_{1})\pi(x_{0},x_{1}),$ (7)
where $p_{0}(x|x_{0},x_{1})=\delta_{x_{0}}(x)$, $p_{1}(x|x_{0},x_{1})=\delta_{x_{1}}(x)$ and $\pi(X_{0},X_{1})$ is an arbitrary joint distribution of $X_{0}$, $X_{1}$ satisfying the marginals constraints $p(x)=\sum_{y\in\mathcal{D}}\pi(x,y)$, $q(y)=\sum_{x\in\mathcal{D}}\pi(x,y)$. Since the probability velocities update the probability independently for each position, it is natural to define $p_{t}(x|x_{0},x_{1})$ independently for each dimension as in Gat et al. (2024):
$p_{t}(x|x_{0},x_{1})=\prod_{i=1}^{N}p^{i}_{t}(x^{i}|x^{i}_{0},x^{i}_{1}),$ (8)
where
$p_{t}^{i}(x^{i}|x_{0},x_{1})=(1-\kappa_{t})\delta_{x_{0}^{i}}(x^{i})+\kappa_{t}\delta_{x_{1}^{i}}(x^{i}),\text{with }\kappa_{0}=0\text{ and }\kappa_{1}=1.$ (9)
It is clear that this definition of $p_{t}(x|x_{0},x_{1})$ satisfies the conditions $p_{0}(x|x_{0},x_{1})=\delta_{x_{0}}(x)$ and $p_{1}(x|x_{0},x_{1})=\delta_{x_{1}}(x)$. In addition, Gat et al. (2024) show that component $i$ of the conditional probability velocity $u_{t}$ corresponding to the flow defined in Equations (8) and (9) is
$u_{t}^{i}(x,z|x_{0},x_{1})=\frac{\dot{\kappa}_{t}}{1-\kappa_{t}}\left[\delta_{x_{1}^{i}}(x^{i})-\delta_{z^{i}}(x^{i})\right].$ (10)
Furthermore, they show that the probability velocity corresponding to the unconditional flow $p_{t}(z)$ can be written as
$u_{t}^{i}(x,z)=\sum_{x_{0},x_{1}\in\mathcal{D}}u_{t}^{i}(x,x_{t}|x_{0},x_{1})p(x_{0},x_{1}|z)dx_{0}dx_{1},$ (11)
which in the case of Equations (9) and (10) implies,
$u_{t}^{i}(x^{i},z)=\frac{\dot{\kappa}_{t}}{1-\kappa_{t}}\left[p_{1|t}(x^{i}|z)-\delta_{z}(x^{i})\right].$ (12)
One then models $u_{t}^{i}(x^{i},z)$ by simply modeling $p^{i}_{1|t}(x^{i}|z)$ with a neural network $p^{i}_{1|t}(x^{i}|z;\theta)$ using the cross entropy loss
$L=-\mathbb{E}_{t\sim U(0,1)}\mathbb{E}_{x_{0},x_{1}\sim\pi(x_{0},x_{1})}\mathbb{E}_{x_{t}\sim p_{t|0}(\cdot|x_{0},x_{1})}\sum_{i=1}^{L}\log p^{i}_{1|t}(x_{1}^{i}|x_{t};\theta).$ (13)
It should be mentioned that in Gat et al. (2024), the definition of $p_{t}^{i}(x^{i}|x_{0},x_{1})$ is given in a more general form, but here we focus on this specific case for the sake of simplicity and since this formulation corresponds to shorts paths conditional flows in the continuous framework, that is $X_{t}=(1-t)X_{0}+tX_{1}$.
2.3.3 Source and Target Distributions
As mentioned, points $X_{0}$ and $X_{1}$ are sampled from a joint distribution $\pi(x,y)$, i.e. $(X_{0},X_{1})\sim\pi(X_{0},X_{1})$, satisfying the marginals constraints $p(x)=\sum_{y\in\mathcal{D}}\pi(x,y)$, $q(y)=\sum_{x\in\mathcal{D}}\pi(x,y)$. As a special case, the training pairs $X_{0}$ and $X_{1}$ can be sampled independently, $(X_{0},X_{1})\sim p(X_{0})q(X_{1})$. Common instantiations of source distribution $p$ are:
(i) adding a special token value often referred to as a mask token, denoted here by $m$, and setting the source distribution to contain only the fully masked sequence, i.e., $(X_{0},X_{1})=((m,\dots,m),X_{1})$.
(ii) using uniform distribution over $\mathcal{D}$, which is equivalent to drawing each $x^{i}$ independently to be some value in $[d]$ with equal probability, denoted $p_{u}(x^{i})$.
One can choose dependent couplings as well, for example, $\pi(x_{0},x_{1})=p(x_{0}|x_{1})q(x_{1})$, which Gat et al. (2024) found improves conditional sampling by partially masking inputs with samples of the form
$(X_{0},X_{1})=(\mathbb{I}\cdot X_{1}+(\mathbf{1}-\mathbb{I})\odot(m,\dots,m),X_{1}),$ (14)
where $X_{1}\sim q(X_{1})$ and $\mathbb{I}\in\{0,1\}^{L}$ is a random variable indicating the conditioning, $\odot$ denotes the entry-wise product, and $\mathbf{1}\in\mathbb{R}^{L}$ is the vector of all ones. They refer to this choice as the conditional coupling (C-coupling).
3 Minibatch Optimal Transport in the Context of Discrete Flow Matching
One of the main objectives of current research in flow matching is to reduce the numbers of steps required for generation. This amounts to simplifying paths/trajectories from the source to the target distribution. An additional benefit to rectifying such paths, is that simplified paths are easier to model by a neural network which empirically has been shown to translate to higher quality models. One can simplify the paths by minimizing the kinetic energy as in the dynamic formulation of optimal transport:
[…]
Such an optimization problem is closely related to minimizing the expected path lengths, but penalizes more heavily the larger norms of the vector field. In addition, according to the Benamou-Brenier theorem (Benamou & Brenier, 2000; Tong et al., 2023), the minimum of Expression (15) is equal to the minimal transportation cost in the Kantorovich formulation
[…]
We point out that minimizing $||v_{t}(x_{t})||^{2}=v_{1,t}^{2}+...+v_{d,t}^{2}$, implies that we are trying to minimize the movements of a particle from the current state. Similarly, in the discrete flow case we wish to minimize the expected outflowing mass $u_{t}^{i}(x^{i},x_{t})$ when $x^{i}\neq x_{t}^{i}$, or in other words maximizing the probability $u_{t}^{i}(x_{t}^{i},x_{t})$ that overall the particles remain at the same place.
Therefore the dynamic formulation in discrete flow matching that we try to minimize is
$\int_{0}^{1}\sum_{x_{t}}p(x_{t})\frac{1}{2}\left[\sum_{i=1}^{L}\left(\sum_{x^{i}\neq x_{t}^{i}}u_{t}^{i}(x^{i},x_{t})-u_{t}^{i}(x_{t}^{i},x_{t})\right)\right]dt=\int_{0}^{1}\sum_{x_{t}}p(x_{t})\left[\sum_{i=1}^{L}\sum_{x^{i}\neq x_{t}^{i}}u_{t}^{i}(x^{i},x_{t})\right]dt,$ (17)
We prove that this is equivalent to the Kantorovich formulation in Equation (16), when the cost function $c(x_{0},x_{1})$ is the Hamming distance ($d_{H}$) between two sequences (Corollary 1).
An issue with the categorical dynamic formulation above is that it treats all different tokens equally, while in practice tokens have different degrees of similarity between each other which is reflected in their embeddings. As such, we would like to weight the outflow to each token by some measure of similarity between the donor and recipient of the flow, such that the transition of mass to dissimilar states is particularly penalized. An additional issue is that for large vocabulary sizes, and two sequences sampled from $p(x_{0})$ and $q(x_{1})$ respectively, only a tiny proportion of their positions match exactly. This is especially the case when the source distribution does not contain information about the target data distribution. This implies that we should not expect substantial improvements by optimizing the Kantorovich expression using minibatches.
For these reasons we define the categorical dynamic objective more generally as follows:
$\int_{0}^{1}\sum_{x_{t}}p(x_{t})\left[\sum_{i=1}^{L}\sum_{x^{i}}u_{t}^{i}(x^{i},x_{t})s(x^{i},x_{t}^{i})\right]dt=\int_{0}^{1}\sum_{x_{t}}p(x_{t})\left[\sum_{i=1}^{L}\sum_{x^{i}\neq x_{t}^{i}}u_{t}^{i}(x^{i},x_{t})s(x^{i},x_{t}^{i})\right]dt,$ (18)
where $s(x^{i},x_{t}^{i})\geq 0$ is a similarity metric between two tokens $x^{i}$ and $x_{t}^{i}$, and thus $s(z,z)=0$. Such a similarity metric serves as a penalizing weight for the components of the outflow. In our previous formulation in Equation (17), we used $s(x^{i},x_{t}^{i})=1-\delta_{x_{t}^{i}}(x^{i})$, that is, the discrete metric. Another natural choice is $s(x^{i},x_{t}^{i})=||e_{m}(x^{i})-e_{m}(x_{t}^{i})||^{2}$, which is the $L_{2}$ norm between the embeddings $e_{m}$ of the tokens. Regardless of the similarity metric used, there exists a corresponding Kantorovich formulation, wherein the cost function can be expressed through the similarity metric.
Theorem 3.1.
[…]
$\sum_{i=1}^{L}s(x_{0}^{i},x_{1}^{i}).$ (19)
We provide a proof in Appendix A. When applying this result to our original categorical dynamic formulation (17) we conclude that the corresponding cost function in the Kantorovich formulation is the Hamming distance $d_{H}$:
Corollary 3.2.
If in Theorem 3.1 we choose $s(x^{i},x_{t}^{i})=1-\delta_{x_{t}^{i}}(x^{i})$ then $c(x_{0},x_{1})=\sum_{i=1}^{L}s(x_{0}^{i},x_{1}^{i})=d_{H}(x_{0},x_{1})$.
Interestingly, if we make the choice $s(x^{i},x_{t}^{i})=||e_{m}(x^{i})-e_{m}(x_{t}^{i})||^{2}$, then the cost function becomes simply the $L_{2}$ norm between the embeddings of the sequences similar to continuous flow matching:
Corollary 3.3.
If in Theorem 3.1 we choose $s(x^{i},x_{t}^{i})=||e_{m}(x^{i})-e_{m}(x_{t}^{i})||^{2}$ then $c(x_{0},x_{1})=\sum_{i=1}^{L}||e_{m}(x_{1}^{i})-e_{m}(x_{0}^{i})||^{2}$ that is $c(x_{0},x_{1})=||e_{m}(x_{1})-e_{m}(x_{0})||^{2}$.
Training Stability
Flow matching with minibatch OT training involves two optimization procedures: choosing the optimal minibatch coupling, and then training the flow model. These training steps influence one-another as the minibatch coupling changes the flow dynamics, while the token embedding updates during flow matching modify optimal minibatch coupling when embedding-based similarity metrics are used. While the fact that, at each training step, the model has only a local view of the flow acts as a stabilizer, we can choose to increase the stability by decoupling the embeddings used in the network and those used in minibatch coupling. One such strategy, for example, could be using a moving average of embeddings during minibatch OT coupling.
Source Distributions
In order to reduce the transportation cost, we would like to reduce the average cost function with respect to pairs of minibatches. This expected cost function is reduced when for each sequence in the data minibatch, there likely is a similar sequence in the minibatch drawn from the source distribution. For this reason, we wish to incorporate information from the data distribution into the source distribution. This can be done in several ways, and we propose two approaches below:
i) Defining the source distribution $p_{0}=p_{o}$, where $p_{o}(x)=\prod_{i=1}^{L}p^{tok}_{o}(x^{i})$, and $p^{tok}$ is the empirical probability distribution of tokens in the training set.
ii) Defining the source distribution $p_{0}=p_{n}$, where $p_{n}(x)$ is the distribution over $K$ means in the training dataset, and the magnitude of the distribution at a particular mean is the normalized cardinality of the cluster belonging to that mean.
4 An Upper Bound for the Perplexity in Discrete Flow Matching
One of the most important metrics that is used to evaluate language models is perplexity. Therefore, it is important to be able to calculate the perplexity over a test set in the discrete flow matching framework, or at least be able to bound this quantity. In what follows, we show that such an upper bound can be derived. To this end, we first provide an upper bound for the KL divergence between the end distributions $\bar{p}_{1}$ and $\bar{q}_{1}$ of two different flows $\bar{p}_{t}$ and $\bar{q}_{t}$ with corresponding velocity probabilities $v_{t}(x^{i},x_{t})$ and $w_{t}(x^{i},x_{t})$, where time moves form 00 to $1$. In order to derive such a bound, we adopt the approach of Opper & Sanguinetti (2007, Equation 3) motivated by the fact that discrete flows are Markov chains, as due to their formulation, the probability over states in the next time point depends solely on the current position.
Theorem 4.1.
For two discrete flows $\bar{p}_{t}$ and $\bar{q}_{t}$ with corresponding velocity probabilities $v_{t}(x^{i},x_{t})$ and $w_{t}(x^{i},x_{t})$, the following inequality holds
$D_{KL}(\bar{q}_{1}||\bar{p}_{1})\leq\int_{0}^{1}\sum_{x_{t}}\bar{q}_{t}(x_{t})\sum_{i=1}^{L}\sum_{x^{i}\neq x_{t}^{i}}\left(w_{t}^{i}(x^{i},x_{t})\log{\frac{w_{t}^{i}(x^{i},x_{t})}{v_{t}^{i}(x^{i},x_{t})}}+v_{t}^{i}(x^{i},x_{t})-w_{t}^{i}(x^{i},x_{t})\right)dt+D_{KL}(\bar{q_{0}}||\bar{p_{0}}).$ (20)
[…]
We notice that since $\bar{q}_{0}(x)=p_{0}(x)$ and $\bar{p}_{0}(x)=p_{0}(x)$, then $D_{KL}(\bar{q}_{0}||\bar{p}_{0})=0$. Furthermore $D_{KL}(\bar{q}_{1}(x)||\bar{p}_{1}(x))=D_{KL}(\delta_{x_{1}}(x)||p_{1}(x;\theta))=-\log{p_{1}(x_{1};\theta)}$. Thus for such particular choices one gets
$-\log{p_{t}(x_{1};\theta)}\leq\int_{0}^{1}\sum_{x_{t}}p_{t|1}(x_{t}|x_{1})\sum_{i=1}^{L}\sum_{x^{i}\neq x_{t}^{i}}\left(w_{t}^{i}(x^{i},x_{t})\log{\frac{w_{t}^{i}(x^{i},x_{t})}{u_{t}^{i}(x^{i},x_{t};\theta)}}+u_{t}^{i}(x^{i},x_{t};\theta)-w_{t}^{i}(x^{i},x_{t})\right)dt.$ (21)
Now we return to the special discrete flow matching dynamics from Equation (9). The probability velocity for $\bar{p}_{t}(x)=p_{t}(x;\theta)$ is given in Equation (10), with the learned velocity being $u_{t}^{i}(x^{i},x_{t};\theta)=\frac{\dot{\kappa}_{t}}{1-\kappa_{t}}\left[p_{1|t}(x^{i}|x_{t};\theta)-\delta_{x_{t}}(x^{i})\right]$. The probability velocity for $\bar{q}_{t}(x)=p_{t|1}(x|x_{1})$ can be calculated by first calculating $p_{t|1}^{i}(x|x_{1})$ using $p_{t|1,0}^{i}(x|x_{1},x_{0})$ in Equation (9), and finding its probability velocity. However this is not necessary because we notice that for $p_{t|1,0}^{i}(x^{i}|x_{1},x_{0})=(1-k_{t})\delta_{x_{0}^{i}}(x^{i})+k_{t}\delta_{x_{1}^{i}}(x^{i})$ the corresponding probability velocity is $u_{t}^{i}(x^{i},x_{t},|x_{0},x_{1})=\frac{\dot{k_{t}}}{1-k_{t}}[\delta_{x_{1}^{i}}(x^{i})-\delta_{x_{t}^{i}}(x^{i})]$ which does not depend on $x_{0}$, thus $w_{t}^{i}(x^{i},x_{t})=u_{t}^{i}(x^{i},x_{t},|x_{1})=u_{t}^{i}(x^{i},x_{t},|x_{0},x_{1})=\frac{\dot{k_{t}}}{1-k_{t}}[\delta_{x_{1}^{i}}(x^{i})-\delta_{x_{t}^{i}}(x^{i})]$. Plugging everything into Expression (21), we get
$-\log{p_{1}(x_{1};\theta)}\leq\int_{0}^{1}\frac{\dot{k_{t}}}{1-k_{t}}\sum_{x_{t}}p_{t|1}(x_{t}|x_{1})\sum_{i=1}^{L}\left(-\delta_{x_{1}^{i}\neq x_{t}^{i}}\log{p_{1|t}^{i}(x^{i}_{1}|x_{t};\theta)}+1-p_{1|t}^{i}(x_{t}^{i}|x_{t};\theta)-\delta_{x_{1}^{i}\neq x_{t}^{i}}\right)dt,$ (22)
therefore, taking the expectation with respect to $p_{1}(x_{1})$, we find that the cross entropy is bounded as follows
$H(p_{1}(x_{1}),p_{1}(x_{1};\theta))\leq\int_{0}^{1}\frac{\dot{k_{t}}}{1-k_{t}}\sum_{x_{t},x_{1}}p_{t,1}(x_{t},x_{1})\sum_{i=1}^{L}\left(-\delta_{x_{1}^{i}\neq x_{t}^{i}}\log{p_{1|t}^{i}(x^{i}_{1}|x_{t};\theta)}+1-p_{1|t}^{i}(x_{t}^{i}|x_{t};\theta)-\delta_{x_{1}^{i}\neq x_{t}^{i}}\right)dt.$ (23)
Finally, since the part inside the large brackets is not dependent on $x_{0}$, we can write
$\mathcal{B}=\int_{0}^{1}\frac{\dot{k_{t}}}{1-k_{t}}\sum_{x_{1},x_{0}}\pi(x_{1},x_{0})\sum_{x_{t}}p_{t|1,0}(x_{t}|x_{1},x_{0})\sum_{i=1}^{L}\left(-\delta_{x_{1}^{i}\neq x_{t}^{i}}\log{p_{1|t}^{i}(x^{i}_{1}|x_{t};\theta)}+1-p_{1|t}^{i}(x_{t}^{i}|x_{t};\theta)-\delta_{x_{1}^{i}\neq x_{t}^{i}}\right)dt,$ (24)
hence $e^{\frac{\mathcal{B}}{L}}$ is a computable upper bound of the perplexity.
5 Future Work: Chains of Discrete Flows
Instead of trying to use a single flow to transform the source distribution into the data distribution, we can be less ambitions and define several flows to perform this task, each refining the modeled distribution of the previous flow. That is, we can define flows $p_{t}^{1},...,p_{t}^{k}$ such that the source distribution $p_{0}$ is transformed successively into $p^{1}_{1},...,p^{k}_{1}\approx p_{data}$. This approach should be particularly beneficial when utilizing minibatch OT for the following reasons. When we try to learn an optimal coupling for the given minibatches, since one of the minibatches is sampled from an unstructured source distribution, the expected similarity between the samples from the target and the source distribution should be low. However, if we use an approximation of data distribution as our source, then we are more likely to find sequences in this distribution that are similar to the ones from the data distribution. However, for example in the case of a chain of two flows, such an approach would require training an additional network, and furthermore it would require using the first network for generation during the training of the second network. As such it remains to be tested if small models with few generation steps in the first flow can substantially increase the quality of the second flow, such that the overall compute/memory per performance ratio is decreased.
6 Conclusion
In this work, we have provided a minimization objective that can be seen as a categorical dynamic formulation of optimal transport to accommodate the unique challenges posed by discrete flows, particularly when addressing the non-deterministic nature of discrete paths which disrupts the application of rectification strategies. By deriving the equivalent Kantorovich formulation, we have enabled the minimization of transitions to dissimilar states by optimizing the transport cost through a minibatch strategy.
Furthermore, we introduced an upper bound on the perplexity for discrete flow models, enabling performance evaluation and comparison to other modeling strategies. These contributions advance our capability to model categorical data with discrete flow models, offering new avenues for efficient and effective data distribution modeling.
Appendix A Theoretical Results
Proof of Theorem 3.1:
We begin with
$\int_{0}^{1}\sum_{x_{t}}p(x_{t})\left(\sum_{i=1}^{L}\sum_{x^{i}\neq x_{t}^{i}}u_{t}^{i}(x^{i},x_{t})s(x^{i},x_{t}^{i})\right)dt$ (25)
which due to Equation (11) can be rewritten as
$\int_{0}^{1}\sum_{x_{t}}p(x_{t})\left(\sum_{i=1}^{L}\sum_{x^{i}\neq x_{t}^{i}}\sum_{x_{0},x_{1}}u_{t}^{i}(x^{i},x_{t}|x_{0},x_{1})p(x_{0},x_{1}|x_{t})s(x^{i},x_{t}^{i})\right)dt=$ (26)
$\int_{0}^{1}\sum_{x_{t}}\sum_{i=1}^{L}\sum_{x^{i}\neq x_{t}^{i}}\sum_{x_{0},x_{1}}u_{t}^{i}(x^{i},x_{t}|x_{0},x_{1})s(x^{i},x_{t}^{i})p(x_{0},x_{1},x_{t})dt=$ (27)
$\int_{0}^{1}\sum_{x_{0},x_{1}}\sum_{i=1}^{L}\sum_{x_{t}}\sum_{x^{i}\neq x_{t}^{i}}u_{t}^{i}(x^{i},x_{t}|x_{0},x_{1})s(x^{i},x_{t}^{i})p(x_{t}|x_{0},x_{1})p(x_{0},x_{1})dt.$ (28)
For $p(x_{t}|x_{0},x_{1})$ as in Equation (9), by Equation (10) we have that
$u_{t}^{i}(x^{i},x_{t}|x_{0},x_{1})=\frac{\dot{k}_{t}}{1-k_{t}}\left(\delta_{x_{1}^{i}}(x^{i})-\delta_{x_{t}^{i}}(x^{i})\right).$ (29)
Continuing from Equation (28)
$\int_{0}^{1}\sum_{x_{0},x_{1}}\sum_{i=1}^{L}\sum_{x_{t}}\sum_{x^{i}\neq x_{t}^{i}}\frac{\dot{k}_{t}}{1-k_{t}}\left(\delta_{x_{1}^{i}}(x^{i})-\delta_{x_{t}^{i}}(x^{i})\right)s(x^{i},x_{t}^{i})p(x_{t}|x_{0},x_{1})p(x_{0},x_{1})dt=$ (30)
$\int_{0}^{1}\sum_{x_{0},x_{1}}\sum_{i=1}^{L}\sum_{x_{t}^{i}}\sum_{x^{i}\neq x_{t}^{i}}\frac{\dot{k}_{t}}{1-k_{t}}\left(\delta_{x_{1}^{i}}(x^{i})-\delta_{x_{t}^{i}}(x^{i})\right)s(x^{i},x_{t}^{i})p(x_{t}^{i}|x_{0},x_{1})p(x_{0},x_{1})dt=$ (31)
$\int_{0}^{1}\sum_{x_{0},x_{1}}\sum_{i=1}^{L}\sum_{\begin{subarray}{c}x_{t}^{i},x^{i}\\ x_{t}^{i}\neq x^{i}\end{subarray}}\frac{\dot{k}_{t}}{1-k_{t}}\left(\delta_{x_{1}^{i}}(x^{i})-\delta_{x_{t}^{i}}(x^{i})\right)s(x^{i},x_{t}^{i})p(x_{t}^{i}|x_{0},x_{1})p(x_{0},x_{1})dt=$ (32)
$\int_{0}^{1}\sum_{x_{0},x_{1}}\sum_{i=1}^{L}\sum_{\begin{subarray}{c}x_{t}^{i},x^{i}\\ x_{t}^{i}\neq x^{i}\end{subarray}}\frac{\dot{k}_{t}}{1-k_{t}}\delta_{x_{1}^{i}}(x^{i})s(x^{i},x_{t}^{i})p(x_{t}^{i}|x_{0},x_{1})p(x_{0},x_{1})dt=$ (33)
$\int_{0}^{1}\sum_{x_{0},x_{1}}\sum_{i=1}^{L}\sum_{\begin{subarray}{c}x_{t}^{i},x^{i}\\ x_{t}^{i}\neq x^{i}\\ x^{i}=x_{1}^{i}\end{subarray}}\frac{\dot{k}_{t}}{1-k_{t}}s(x^{i},x_{t}^{i})p(x_{t}^{i}|x_{0},x_{1})p(x_{0},x_{1})dt$ (34)
where again due to the choice of Equation (9)
$\int_{0}^{1}\sum_{x_{0},x_{1}}\sum_{i=1}^{L}\sum_{\begin{subarray}{c}x_{t}^{i},x^{i}\\ x_{t}^{i}\neq x^{i}\\ x^{i}=x_{1}^{i}\end{subarray}}\frac{\dot{k}_{t}}{1-k_{t}}s(x^{i},x_{t}^{i})\left((1-k_{t})\delta_{x_{0}^{i}}(x_{t}^{i})+k_{t}\delta_{x_{1}^{i}}(x_{t}^{i})\right)p(x_{0},x_{1})dt=$ (35)
$\int_{0}^{1}\sum_{x_{0},x_{1}}\sum_{i=1}^{L}\left(\sum_{\begin{subarray}{c}x_{t}^{i},x^{i}\\ x_{t}^{i}\neq x^{i}\\ x^{i}=x_{1}^{i}\end{subarray}}\frac{\dot{k}_{t}}{1-k_{t}}(1-k_{t})\delta_{x_{0}^{i}}(x_{t}^{i})+\sum_{\begin{subarray}{c}x_{t}^{i},x^{i}\\ x_{t}^{i}\neq x^{i}\\ x^{i}=x_{1}^{i}\end{subarray}}\frac{\dot{k}_{t}}{1-k_{t}}k_{t}\delta_{x_{1}^{i}}(x_{t}^{i})\right)s(x^{i},x_{t}^{i})p(x_{0},x_{1})dt=$ (36)
$\int_{0}^{1}\sum_{x_{0},x_{1}}\sum_{i=1}^{L}\left(\sum_{\begin{subarray}{c}x_{t}^{i},x^{i}\\ x_{t}^{i}\neq x^{i}\\ x^{i}=x_{1}^{i}\end{subarray}}\dot{k}_{t}\delta_{x_{0}^{i}}(x_{t}^{i})+0\right)s(x^{i},x_{t}^{i})p(x_{0},x_{1})dt$ (37)
where the second expression is zero since in the sum one must have $x_{t}^{i}\neq x^{i}$ and $x^{i}=x_{1}^{i}$, therefore $x_{t}^{i}\neq x_{1}^{i}$ which sets $\delta_{x_{1}^{i}}(x_{t}^{i})$ to 0. Hence we only have
$\int_{0}^{1}\sum_{x_{0},x_{1}}\sum_{i=1}^{L}\sum_{\begin{subarray}{c}x_{t}^{i},x^{i}\\ x_{t}^{i}\neq x^{i}\\ x^{i}=x_{1}^{i}\end{subarray}}\dot{k}_{t}\delta_{x_{0}^{i}}(x_{t}^{i})s(x^{i},x_{t}^{i})p(x_{0},x_{1})dt=\int_{0}^{1}\sum_{x_{0},x_{1}}\sum_{i=1}^{L}\sum_{\begin{subarray}{c}x_{t}^{i},x^{i}\\ x_{t}^{i}\neq x^{i}\\ x^{i}=x_{1}^{i}\\ x_{t}^{i}=x_{0}^{i}\end{subarray}}s(x^{i},x_{t}^{i})\dot{k}_{t}p(x_{0},x_{1})dt$ (38)
Expression
$\sum_{\begin{subarray}{c}x_{t}^{i},x^{i}\\ x_{t}^{i}\neq x^{i}\\ x^{i}=x_{1}^{i}\\ x_{t}^{i}=x_{0}^{i}\end{subarray}}s(x^{i},x_{t}^{i})$ (39)
is clearly $s(x_{1}^{i},x_{0}^{i})$ when $x_{0}^{i}\neq x_{1}^{i}$ and zero otherwise. Thus, we have
$\int_{0}^{1}\sum_{x_{t}}p(x_{t})\left(\sum_{i=1}^{L}\sum_{x^{i}\neq x_{t}^{i}}s(x^{i},x_{t}^{i})u_{t}^{i}(x^{i},x_{t})\right)dt=\int_{0}^{1}\sum_{x_{0},x_{1}}\sum_{i=1}^{L}s(x_{1}^{i},x_{0}^{i})\dot{k}_{t}p(x_{0},x_{1})dt=$ (40)
$\sum_{x_{0},x_{1}}\sum_{i=1}^{L}s(x_{1}^{i},x_{0}^{i})(k_{1}-k_{0})p(x_{0},x_{1})=\sum_{x_{0},x_{1}}\sum_{i=1}^{L}s(x_{1}^{i},x_{0}^{i})(1-0)p(x_{0},x_{1})$ (41)
We conclude that
$\int_{0}^{1}\sum_{x_{t}}p(x_{t})\left(\sum_{i=1}^{L}\sum_{x^{i}\neq x_{t}^{i}}u_{t}^{i}(x^{i},x_{t})s(x^{i},x_{t}^{i})\right)dt=\sum_{x_{0},x_{1}}c(x_{0},x_{1})p(x_{0},x_{1}),$ (42)
where $c(x_{0},x_{1})=\sum_{i=1}^{L}s(x_{1}^{i},x_{0}^{i})$.
Proof of Theorem 4.1:
We begin by defining two discrete Markov chains $\hat{p}_{t}$ and $\hat{q}_{t}$ whose timestep sizes are $\epsilon$ and the total number of steps is $K=\lfloor\frac{1}{\epsilon}\rfloor$, such that when $\epsilon\rightarrow 0$, their marginal distributions converge to those of the flows $\bar{p}_{t}$ and $\bar{q}_{t}$. The KL divergence between the paths of such Markov chains can be written as below:
$D_{KL}(\hat{q},\hat{p})=\sum_{x_{0:K\epsilon}}\hat{q}(x_{0:K\epsilon})\log{\frac{\hat{q}(x_{0:K\epsilon})}{\hat{p}(x_{0:K\epsilon})}}=\sum_{x_{0:K\epsilon}}\hat{q}(x_{0:K\epsilon})\log{\prod_{k=1}^{K}\frac{\hat{q}(x_{k\epsilon}|x_{(k-1)\epsilon},...,x_{0})}{\hat{p}(x_{k\epsilon}|x_{(k-1)\epsilon},...,x_{0})}\frac{\hat{q}(x_{0})}{\hat{p}(x_{0})}}$ (43)
$=\sum_{x_{0:K\epsilon}}\hat{q}(x_{0:K\epsilon})\left(\sum_{k=1}^{K}\log{\frac{\hat{q}(x_{k\epsilon}|x_{(k-1)\epsilon})}{\hat{p}(x_{k\epsilon}|x_{(k-1)\epsilon})}}+\log{\frac{\hat{q}(x_{0})}{\hat{p}(x_{0})}}\right)$ (44)
$=\sum_{k=1}^{K}\sum_{\begin{subarray}{c}x_{k\epsilon}\\ x_{(k-1)\epsilon}\end{subarray}}\hat{q}(x_{k\epsilon},x_{(k-1)\epsilon})\log{\frac{\hat{q}(x_{k\epsilon}|x_{(k-1)\epsilon})}{\hat{p}(x_{k\epsilon}|x_{(k-1)\epsilon})}}+\sum_{x_{0}}\hat{q}(x_{0})\log{\frac{\hat{q}(x_{0})}{\hat{p}(x_{0})}}$ (45)
$=\sum_{k=1}^{K}\sum_{x_{(k-1)\epsilon}}\hat{q}(x_{(k-1)\epsilon})\sum_{x_{k\epsilon}}\hat{q}(x_{k\epsilon}|x_{(k-1)\epsilon})\log{\frac{\hat{q}(x_{k\epsilon}|x_{(k-1)\epsilon})}{\hat{p}(x_{k\epsilon}|x_{(k-1)\epsilon})}}+\sum_{x_{0}}\hat{q}(x_{0})\log{\frac{\hat{q}(x_{0})}{\hat{p}(x_{0})}}$ (46)
$=I+D_{KL}(\hat{q}(x_{0})||\hat{p}(x_{0})),$ (47)
where
$I=\sum_{k=1}^{K}\sum_{x_{(k-1)\epsilon}}\hat{q}(x_{(k-1)\epsilon})\sum_{x_{k\epsilon}}\hat{q}(x_{k\epsilon}|x_{(k-1)\epsilon})\log{\frac{\hat{q}(x_{k\epsilon}|x_{(k-1)\epsilon})}{\hat{p}(x_{k\epsilon}|x_{(k-1)\epsilon})}}$ (48)
is a weighted sum of KL divergences with non-negative weights, that is
$I=\sum_{k=1}^{K}\sum_{x_{(k-1)\epsilon}}\hat{q}(x_{(k-1)\epsilon})D_{KL}(\hat{q}(x_{k\epsilon}|x_{(k-1)\epsilon})||\hat{p}(x_{k\epsilon}|x_{(k-1)\epsilon})).$ (49)
Now, we focus on computing expression $D_{KL}(\hat{q}(x_{k\epsilon}|x_{(k-1)\epsilon})||\hat{p}(x_{k\epsilon}|x_{(k-1)\epsilon}))$. The sum
$\sum_{x_{k\epsilon}}\hat{q}(x_{k\epsilon}|x_{(k-1)\epsilon})\log{\frac{\hat{q}(x_{k\epsilon}|x_{(k-1)\epsilon})}{\hat{p}(x_{k\epsilon}|x_{(k-1)\epsilon})}}$ (50)
can be separated into three sums:
$\sum_{\begin{subarray}{c}x_{k\epsilon}\\ d_{H}(x_{k\epsilon},x_{(k-1)\epsilon})=0\end{subarray}}\hat{q}(x_{k\epsilon}|x_{(k-1)\epsilon})\log{\frac{\hat{q}(x_{k\epsilon}|x_{(k-1)\epsilon})}{\hat{p}(x_{k\epsilon}|x_{(k-1)\epsilon})}}+\sum_{\begin{subarray}{c}x_{k\epsilon}\\ d_{H}(x_{k\epsilon},x_{(k-1)\epsilon})=1\end{subarray}}\hat{q}(x_{k\epsilon}|x_{(k-1)\epsilon})\log{\frac{\hat{q}(x_{k\epsilon}|x_{(k-1)\epsilon})}{\hat{p}(x_{k\epsilon}|x_{(k-1)\epsilon})}}$ (51)
$+\sum_{\begin{subarray}{c}x_{k\epsilon}\\ d_{H}(x_{k\epsilon},x_{(k-1)\epsilon})>1\end{subarray}}\hat{q}(x_{k\epsilon}|x_{(k-1)\epsilon})\log{\frac{\hat{q}(x_{k\epsilon}|x_{(k-1)\epsilon})}{\hat{p}(x_{k\epsilon}|x_{(k-1)\epsilon})}}.$ (52)
We first analyze the second sum. Since $x_{k\epsilon}$ and $x_{(k-1)\epsilon}$ differ at exactly one neighbour (say position $j$), from the flow matching update rule $p^{i}_{k\epsilon|(k-1)\epsilon}(x^{i}|x_{(k-1)\epsilon})=\delta_{x_{(k-1)\epsilon}^{i}}(x^{i})+\epsilon u^{i}(x^{i},x_{(k-1)\epsilon})$ applied independently to each position, we can infer that
$p_{k\epsilon|(k-1)\epsilon}(x_{k\epsilon}|x_{(k-1)\epsilon})=\epsilon u^{j}(x^{j},x_{(k-1)\epsilon})\prod_{\begin{subarray}{c}i=1\\ i\neq j\end{subarray}}^{L}\left(1+u^{i}(x^{i},x_{(k-1)\epsilon})\epsilon\right)=\epsilon u^{j}(x^{j},x_{(k-1)\epsilon})+O(\epsilon^{2})$ (53)
therefore
$\sum_{\begin{subarray}{c}x_{k\epsilon}\\ d_{H}(x_{k\epsilon},x_{(k-1)\epsilon})=1\end{subarray}}\hat{q}(x_{k\epsilon}|x_{(k-1)\epsilon})\log{\frac{\hat{q}(x_{k\epsilon}|x_{(k-1)\epsilon})}{\hat{p}(x_{k\epsilon}|x_{(k-1)\epsilon})}}=$ (54)
$=\sum_{j=1}^{L}\sum_{x^{j}\neq x^{j}_{(k-1)\epsilon}}\epsilon w^{j}(x^{j},x_{(k-1)\epsilon})\log{\frac{w^{j}(x^{j},x_{(k-1)\epsilon})+O(\epsilon)}{v^{j}(x^{j},x_{(k-1)\epsilon})+O(\epsilon)}}+O(\epsilon^{2}).$ (55)
For the third sum, since
$p_{k\epsilon|(k-1)\epsilon}(x_{k\epsilon}|x_{(k-1)\epsilon})=\epsilon^{2}u^{j}(x^{j},x_{(k-1)\epsilon})u^{l}(x^{j},x_{(k-1)\epsilon})\prod_{\begin{subarray}{c}i=1\\ i\neq j,l\end{subarray}}^{L}\left(1+u^{i}(x^{i},x_{(k-1)\epsilon})\epsilon\right)=O(\epsilon^{2})+O(\epsilon^{3})=O(\epsilon^{2})$ (56)
we conclude that
$\sum_{\begin{subarray}{c}x_{k\epsilon}\\ d_{H}(x_{k\epsilon},x_{(k-1)\epsilon})>1\end{subarray}}\hat{q}(x_{k\epsilon}|x_{(k-1)\epsilon})\log{\frac{\hat{q}(x_{k\epsilon}|x_{(k-1)\epsilon})}{\hat{p}(x_{k\epsilon}|x_{(k-1)\epsilon})}}=O(\epsilon^{2}).$ (57)
Therefore the only sum left is the first one
$\sum_{\begin{subarray}{c}x_{k\epsilon}\\ d_{H}(x_{k\epsilon},x_{(k-1)\epsilon})=0\end{subarray}}\hat{q}(x_{k\epsilon}|x_{(k-1)\epsilon})\log{\frac{\hat{q}(x_{k\epsilon}|x_{(k-1)\epsilon})}{\hat{p}(x_{k\epsilon}|x_{(k-1)\epsilon})}}=\hat{q}(x_{(k-1)\epsilon}|x_{(k-1)\epsilon})\log{\frac{\hat{q}(x_{(k-1)\epsilon}|x_{(k-1)\epsilon})}{\hat{p}(x_{(k-1)\epsilon}|x_{(k-1)\epsilon})}}.$ (58)
In this special case,
$p_{k\epsilon|(k-1)\epsilon}(x_{(k-1)\epsilon}|x_{(k-1)\epsilon})=\prod_{\begin{subarray}{c}i=1\end{subarray}}^{L}\left(1+u^{i}(x^{i}=x_{(k-1)\epsilon}^{i},x_{(k-1)\epsilon})\epsilon\right)$ (59)
$=1+\epsilon\sum_{i=1}^{L}u^{i}(x^{i}=x_{(k-1)\epsilon}^{i},x_{(k-1)\epsilon})+O(\epsilon^{2}),$ (60)
thus
$\hat{q}(x_{(k-1)\epsilon}|x_{(k-1)\epsilon})=1+\epsilon\sum_{i=1}^{L}w^{i}(x^{i}=x_{(k-1)\epsilon}^{i},x_{(k-1)\epsilon})+O(\epsilon^{2}),$ (61)
$\log{\hat{q}(x_{(k-1)\epsilon}|x_{(k-1)\epsilon})}=\epsilon\sum_{i=1}^{L}w^{i}(x^{i}=x_{(k-1)\epsilon}^{i},x_{(k-1)\epsilon})+O(\epsilon^{2}),$ (62)
$\log{\hat{p}(x_{(k-1)\epsilon}|x_{(k-1)\epsilon})}=\epsilon\sum_{i=1}^{L}v^{i}(x^{i}=x_{(k-1)\epsilon}^{i},x_{(k-1)\epsilon})+O(\epsilon^{2}),$ (63)
implying
$\hat{q}(x_{(k-1)\epsilon}|x_{(k-1)\epsilon})\log{\frac{\hat{q}(x_{(k-1)\epsilon}|x_{(k-1)\epsilon})}{\hat{p}(x_{(k-1)\epsilon}|x_{(k-1)\epsilon})}}=\hat{q}(x_{(k-1)\epsilon}|x_{(k-1)\epsilon})\left(\log{\hat{q}(x_{(k-1)\epsilon}|x_{(k-1)\epsilon})}-\log{\hat{p}(x_{(k-1)\epsilon}|x_{(k-1)\epsilon})}\right)$ (64)
$=\epsilon\sum_{i=1}^{L}w^{i}(x^{i}=x_{(k-1)\epsilon}^{i},x_{(k-1)\epsilon})-\epsilon\sum_{i=1}^{L}v^{i}(x^{i}=x_{(k-1)\epsilon}^{i},x_{(k-1)\epsilon})+O(\epsilon^{2}).$ (65)
When accounting for the fact that $u(x^{i}=x_{t}^{i},x_{t})=-\sum_{x^{i}\neq x_{t}^{i}}u(x^{i},x_{t})$, we finally get
$\hat{q}(x_{(k-1)\epsilon}|x_{(k-1)\epsilon})\log{\frac{\hat{q}(x_{(k-1)\epsilon}|x_{(k-1)\epsilon})}{\hat{p}(x_{(k-1)\epsilon}|x_{(k-1)\epsilon})}}=\epsilon\sum_{i=1}^{L}\sum_{x^{i}\neq x_{(k-1)\epsilon}^{i}}\left(v^{i}(x^{i},x_{(k-1)\epsilon})-w^{i}(x^{i},x_{(k-1)\epsilon})\right)+O(\epsilon^{2}).$ (66)
We get the expression for $D_{KL}(\hat{q}(x_{k\epsilon}|x_{(k-1)\epsilon})||\hat{p}(x_{k\epsilon}|x_{(k-1)\epsilon}))$ by adding all three sums,
$\epsilon\sum_{i=1}^{L}\sum_{x^{i}\neq x_{(k-1)\epsilon}^{i}}\left(w^{i}(x^{i},x_{(k-1)\epsilon})\log{\frac{w^{i}(x^{i},x_{(k-1)\epsilon})+O(\epsilon)}{v^{i}(x^{i},x_{(k-1)\epsilon})+O(\epsilon)}}+v^{i}(x^{i},x_{(k-1)\epsilon})-w^{i}(x^{i},x_{(k-1)\epsilon})\right)+O(\epsilon^{2}).$ (67)
Writing $t_{k}=x_{k-1}\epsilon$, and plugging this last expression in $I$, one gets
$I=\sum_{k=0}^{K-1}\epsilon\sum_{x_{t_{k}}}\hat{q}(x_{t_{k}})\sum_{i=1}^{L}\sum_{x^{i}\neq x_{t_{k}}^{i}}\left(w_{t_{k}}^{i}(x^{i},x_{t_{k}})\log{\frac{w_{t_{k}}^{i}(x^{i},x_{t_{k}})+O(\epsilon)}{v_{t_{k}}^{i}(x^{i},x_{t_{k}})+O(\epsilon)}}+v_{t_{k}}^{i}(x^{i},x_{t_{k}})-w_{t_{k}}^{i}(x^{i},x_{t_{k}})\right)+O(\epsilon).$ (68)
Finally taking the limit $\epsilon\rightarrow 0$
$\int_{0}^{1}\sum_{x_{t}}\bar{q}(x_{t})\sum_{i=1}^{L}\sum_{x^{i}\neq x_{t}^{i}}\left(w_{t}^{i}(x^{i},x_{t})\log{\frac{w_{t}^{i}(x^{i},x_{t})}{v_{t}^{i}(x^{i},x_{t})}}+v_{t}^{i}(x^{i},x_{t})-w_{t}^{i}(x^{i},x_{t})\right)dt.$ (69)
Based on the last formula we can write
$D_{KL}(\bar{q},\bar{p})=\int_{0}^{1}\sum_{x_{t}}\bar{q}_{t}(x_{t})\sum_{i=1}^{L}\sum_{x^{i}\neq x_{t}^{i}}\left(w_{t}^{i}(x^{i},x_{t})\log{\frac{w_{t}^{i}(x^{i},x_{t})}{v_{t}^{i}(x^{i},x_{t})}}+v_{t}^{i}(x^{i},x_{t})-w_{t}^{i}(x^{i},x_{t})\right)dt+D_{KL}(\bar{q}(x_{0})||\bar{p}(x_{0})).$ (70)
We now set to prove that $D_{KL}(\bar{q}(x_{1})||\bar{p}(x_{1}))\leq D_{KL}(\bar{q},\bar{p})$. The proof will be an immediate consequence of the Jensen inequality. Indeed,
$-D_{KL}(\hat{q}||\hat{p})=\sum_{x_{0:K\epsilon}}\hat{q}(x_{0:K\epsilon})\log{\frac{\hat{p}(x_{0:K\epsilon})}{\hat{q}(x_{0:K\epsilon})}}=\sum_{x_{0:K\epsilon}}\hat{q}(x_{0})\hat{q}(x_{\epsilon:K\epsilon}|x_{0})\log{\frac{\hat{p}(x_{0:K\epsilon})}{\hat{q}(x_{0:K\epsilon})}}$ (71)
$=\sum_{x_{0}}\hat{q}(x_{0})\sum_{x_{\epsilon:K\epsilon}}\hat{q}(x_{\epsilon:K\epsilon}|x_{0})\log{\frac{\hat{p}(x_{0:K\epsilon})}{\hat{q}(x_{0:K\epsilon})}}\leq\sum_{x_{0}}\hat{q}(x_{0})\log\sum_{x_{\epsilon:K\epsilon}}\hat{q}(x_{\epsilon:K\epsilon}|x_{0}){\frac{\hat{p}(x_{0:K\epsilon})}{\hat{q}(x_{0:K\epsilon})}}$ (72)
$\sum_{x_{0}}\hat{q}(x_{0})\log\sum_{x_{\epsilon:K\epsilon}}{\frac{\hat{p}(x_{0:K\epsilon})}{\hat{q}(x_{0})}}=-D_{KL}(\hat{q}(x_{0})||\hat{p}(x_{0}))$ (73)
Therefore, $D_{KL}(\hat{q}_{0}||\hat{p}_{0})\leq D_{KL}(\hat{q}||\hat{p})$, and taking the limit $\epsilon\rightarrow 0$ we get $D_{KL}(\bar{q}_{0}||\bar{p}_{0})\leq D_{KL}(\bar{q}||\bar{p})$. Applying this result to the reverse processes that generate the marginals $\bar{p}$ and $\bar{q}$ gives $D_{KL}(\bar{q}_{1}||\bar{p}_{1})\leq D_{KL}(\bar{q},\bar{p})$.
In total we have proved that
$D_{KL}(\bar{q}_{1}||\bar{p}_{1})\leq\int_{0}^{1}\sum_{x_{t}}\bar{q}_{t}(x_{t})\sum_{i=1}^{L}\sum_{x^{i}\neq x_{t}^{i}}\left(w_{t}^{i}(x^{i},x_{t})\log{\frac{w_{t}^{i}(x^{i},x_{t})}{v_{t}^{i}(x^{i},x_{t})}}+v_{t}^{i}(x^{i},x_{t})-w_{t}^{i}(x^{i},x_{t})\right)dt+D_{KL}(\bar{q}_{0}||\bar{p}_{0}).$ (74)